# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ IX.3 Power coupling between the macroscopic system and the microscopic arc development
The above formulation is applicable to any high-field system and geometry, DC or RF. It provides a general and rigorous formalism to connect the macroscopic experimental findings of breakdown rate dependence to the microscopic arc ignition simulation models presented in sections VI and VII. Furthermore, eq. (10) is a generalized version of the external coupling circuit model already implemented in ArcPIC and FEMOCS. In fact, the RC circuit implementation of [166, 80] is a special case of (10).
A systematic simulation of the full coupling of this model with PIC is an ongoing work. However, simple analytic approximations for the impedance functions can be considered for various systems, providing valuable insights and an intuitive understanding of the processes that impede plasma development and result in the empirically observed trends described above. Furthermore, by considering such a simplified approach, as proposed by Paszkiewicz et. al. [136], a direct comparison with empirical data is possible, yielding a very good prediction of the empirically observed limits.
@@
A second simplification that can be done for RF accelerating structures, which gives quite a good agreement with empirical data, is to look at the value of at the single operating frequency of the accelerating structure. At this frequency, the system becomes equivalent to a series resistor circuit, yielding a simple picture of the breakdown impedance, graphically depicted in figure 68.
In this figure, the straight lines represent the linear dependence of on for three cases, with the slope of the line defining the breakdown impedance . The black curve gives the BD emission characteristic. The crossing point corresponds to the operating field during the breakdown, which Paszkiewicz et al. proposed as a new empirical limit. In other words, locations in a structure with a higher operating field are expected to have a higher breakdown rate as they maintain a higher flow of electromagnetic power into the breakdown.
The value of needs to be calculated numerically, but simple models can yield analytical formulae. Such a formula has been derived [136]. by modeling a TW accelerating structure as a series of coupled LCR circuits, each representing a cell of the RF structure. This model yielded
which was also confirmed by numerical simulations. The dependence on group velocity, as well as the ratio of surface to accelerating fields, appears consistent with the early empirical predictors of breakdown performance, such as the group velocity, or , as the field ratio scales with the circumference.
Beyond this simplification, the value of has been calculated numerically for multiple systems, yielding very promising results that agree with experimental observations. These results can be found in ref. [136]. Here, we shall present the results for two specific systems, namely the DC large electrode system and the crab cavity.
Figure 69 demonstrates the success of this approach for the crab cavity, in contrast with the picture of figure 67. The distribution of the loaded field follows the same trend as the unloaded electric field , thus restoring the prediction that the high-field performance is determined by the power flow rather than the surface electric field value.
Another system where previous high field limits (, etc.) failed to predict the breakdown distribution is the LES system, as group velocity and are not definable in DC in general. However, the above approach gave excellent predictions. One of the approximations that is necessary to abandon when studying the DC system is the single frequency. In the DC, there is no reason to choose a certain operating frequency. For that reason, instead of , the indicating limit shall be given by the RMS value of over a wide range of frequencies. Using this RMS average as a metric, both the positional dependence of breakdown density inside the surface of LES systems and the gap scaling have been successfully predicted [136].
@@ X.1 Current understanding
In this review, we have summarised a body of experimental, theoretical, and simulation studies that address the fundamental processes of vacuum breakdown initiation as well as basic questions of how electric field holding by metal surfaces depends on a wide range of parameters and conditions. Several notable results significantly advanced the overall understanding of vacuum breakdown. One important aspect of the advance is that experimental work within the included studies was carried out in such a way that data could be quantitatively compared and combined. The resulting data consequently covers high-field and breakdown phenomena over a much wider range of conditions and system types that have never been assembled and analyzed together before. The collective work contributes both to fundamental understanding and practical considerations for real-world devices.
@@
One of the most striking theoretical insights coming out of the work is the identification of the central role in the breakdown initiation of the plastic response, i.e., the dislocation dynamics of the electrode material to the stresses caused by surface electromagnetic fields. This insight explains the ordering of field-holding strength of different materials observed in pulsed DC experiments through the different material crystal structures and corresponding energies required for dislocation motion. A formalism based on dislocation dynamics provides a predictive model for the observed dependence of breakdown rate on the applied electric field in both DC and RF devices. This very strong dependence has been observed and measured in multiple RF structures, structure types, and pulsed DC electrodes. The same formalism also accurately predicts the observed dependence of field-holding on temperature, particularly the striking increase in field-holding as high-field surfaces are cooled to cryogenic temperatures. In addition, the dislocation model provides explanations for details of experimentally observed breakdown behavior, for example, the pulse-to-pulse memory effect that is deduced from the experimental result that the distribution of the timing of breakdowns within pulses is uniform, even though there is a strong dependence of breakdown rate on pulse length.
@@
Analysis of systematic and detailed data of the conditioning and long-term operation from the many tens of high-field RF and DC systems covered in this review shows that the underlying conditioning process correlates to the number of pulses, rather than the number of breakdowns. This observation indicates the existence of intrinsic features that can condition under cyclic tensile stress exerted on metal surfaces by applied electric fields during the high-voltage pulses. The conditioning of intrinsic features occurs in addition to the conditioning of extrinsic features that dominate the conditioning of lower electric field level systems covered by previous studies. The former dominates at high electric fields. The way in which the field holding capability of a device evolves in response to the application of fields has practical importance for operational devices but also gives strong insight into the breakdown initiation process. The strong correlation to the number of pulses gives strong evidence that the evolution of the dislocation structure is responsible for the long-term conditioning of high-field devices. This insight into the pulse number dependence has allowed the optimization of conditioning strategies, aided by a Monte Carlo simulation tool that models the conditioning and breakdown process.
@@
Data from testing of multiple RF structures with different RF designs, and consequently RF properties, show a remarkable dependence of the peak surface field holding capacity, a local quantity, on the overall RF structure geometry. This means that the macroscopic geometry of the RF structures has an influence on the microscopic breakdown initiation process. The key element in explaining this macroscopic to microscopic influence has been to identify the central role of power flow. The simple statement is that an inflow of power is necessary to support the breakdown process, and in particular when the plasma is forming and ever-increasing currents are emitted from the breakdown site. A series of ever more complete theoretical models based on the coupling of power flow to breakdowns has been developed. The earliest models considered global power flow, a later one local complex power flow and finally a full dynamic beam-loading model has been developed. These successive generations of models quantitatively predict an ever wider range of experiments, RF and also DC electrodes, the latter through gap dependence. The unprecedented availability of high-quality, quantitatively comparable high-field data from a wide range of devices from the linear collider development programs was necessary to guide these theoretical developments.
In addition to dislocation dynamics under the metal surface, the dynamics of surface atoms has also been taken into account. Although dislocation activities triggered by applied electric fields may result in surface irregularities, the latter are usually fairly small and blunt. These features alone are not able to explain the strong increase of electric fields measured before a breakdown event takes place. The modeling of surface dynamics using the tools developed specifically for this purpose revealed the possibility of biased diffusion of surface atoms that become partially charged and respond to the presence of surface gradient, similarly to the polarised neutral atoms. This process is proposed to contribute to the sharpening and further growth of a surface irregularity triggered at the surface by a dislocation reaction that could have occurred beneath the surface. This process is also thermally activated and hence, slows down at cryogenic temperatures. Moreover, the process is self-reinforcing, the higher and sharper the feature becomes, the stronger the field gradient, hence the faster the growth of the feature proceeds.
Reduction of work function due to the application of an electric field was also considered. It was found unlikely to expect a significant effect due to electron density redistribution around the surface atoms exposed to the external electric field, however, some experimental works report the observed reduction of work function. Some variation of work function value is observed theoretically and experimentally for different surface facets, however, further theoretical studies are needed to investigate whether the bulk defects developed under the applied pulsed electric field, may contribute to the change of the value of the work function dramatically.
@@
A long-standing mystery in data from high-field systems is the nature of the features that cause emission enhancement in macroscopic systems. Some mechanism is needed to reconcile the current vs. applied field measurements with electron field emission theory as expressed by the Fowler-Nordheim equation. A widespread assumption is that there are microscopic surface features, commonly called tips, that cause a geometrical field enhancement which reconciles experiment with theory. Field enhancement factors of 30 are typically needed for the fully conditioned structures described in this report, and factors of 100 to 200 are needed for the DC electrodes. Advanced modeling capabilities have been developed and applied to tips that would give such field enhancements. The models include the response of the tips to the applied field including heating and cooling processes due to current flow and electron emission, mechanical stresses, and emission of copper atoms then also include plasma formation and evolution. This models the evolution process including all important physical processes of the breakdown from an initial condition through the observed transition of the vacuum to a conductive medium and collapse of applied fields. The tip simulations provide a simulation basis for the time evolving impedance used in the power coupling modeling that gives geometrical dependence field holding. The simulations provide insight into different aspects of breakdown sites, including for example constraints on the nature of the tips for example on thermal coupling to base.
@@ X.2 Prospects and open questions
This report provides a coherent theoretical foundation to a large body of mature experimental work. There are, inevitably a number of open questions that are the subject of ongoing efforts, in particular in connecting remaining important theoretical constructs with direct experimental observations. These are being actively pursued at the time of the writing of this report.
@@
One of these ongoing research focus areas is to determine directly the experimental signature of conditioning, that is to identify how a conditioned surface, or subsurface, differs from an unconditioned one. As we have seen conditioning is both of significant practical importance and inherently addresses the question of breakdown nucleation, so this is a very important question. One of the main experimental lines of this effort is to make microscopic imaging of devices at progressive conditioning stages and in complement as a function of position in spatially varying surface fields. The primary experimental tool is electron microscopy, for example with FIB sample preparation followed by TEM analysis. The goal of this analysis is to determine if a change in dislocation structure or population can be identified between conditioned and unconditioned surfaces.
@@
Another ongoing experimental effort is to determine the nature of the enhanced emission sites, and how they populate a high-field surface. At the time of this writing, no sufficiently sharp tip-like feature has ever been found in many hundreds of SEM images taken in the course of the experiments described in this report. This could be due to the small probability of seeing a few-10s of nanometer expected feature size on cm-scale surfaces. Localizing emitter sites is being addressed by imaging of field emission, in particular by upgrading the breakdown positioning cameras of pulsed DC systems to higher sensitivity. This will allow the question of whether breakdowns preferentially occur where stable field emitters are observed, or not, to be answered.
A possible upgrade of the current analysis method which is based mostly on post-mortem microscopy, would be an in-vacuum high-field device, combined with microscopy analysis and possibly with field emission measurements (field emission microscopy and possibly spectroscopy) which may allow in-situ estimates of field-plasticity relationship. Initial attempts at such experiments have been conducted in the past, but the technical challenges have not allowed for robust and conclusive results. On the other hand, the demonstrated plastic evolution is likely occurring at scales that require scanning tunneling microscopy but while a system designed to provide direct access from field exposure to SEM microscopy is possible, sub-surface microscopy during high field exposure is impossible. This would require other independent methods of characterization, allowing access to monitoring plastic response in synchro with the field exposure, such as monitoring radiation, acoustic emission due to dislocation motion, and changes in high-frequency impedance.
Such future experimental independent measurements of sub-threshold plastic activity as well as in-situ constraints are expected to enable the advancement of theoretical models to provide further details of the break-down nucleation event. Such constraints are needed, for example for allowing effective models such as the MDDF to move from 0-dimensional mean-field description to addressing spatial local interactions between plastic sites, which may be of critical importance in deriving a full picture of the nucleation event.
Another open question, the answer of which may have profound consequences both on the theoretical understanding of the physics of arc development, and the practical design of high-field devices such as high-gradient accelerators, is the dependence of the breakdown rates and maximum allowable fields on the electromagnetic characteristics of a device, as described in section IX. Finding a quantity that can reliably predict this dependence is found, it might enable an advanced optimization of the device geometry and electromagnetic parameters to minimize breakdown rates and significantly increase the maximum allowed fields.
The results, methods, and insights are being applied to an increasingly large range of applications and systems. The new applications also often imply additional elements that may enhance breakdown likelihood. One example has been the investigation of possible enhanced breakdown due to H- irradiation in the CERN LINAC4 RFQ, which is part of the LHC injector chain. H- irradiation can cause carbon layers but also creates dislocation and in the extreme case blisters. Testing uses the pulsed DC system. The study of a possible Muon collider aims to understand the role of an externally applied magnetic field. FCC injection kickers would like to understand the role of synchrotron radiation-produced X-rays on breakdown. As a consequence, this body of research work is being actively carried forward in an expanding community.

### v2

@@ IX.3 Power coupling between the macroscopic system and the microscopic arc development
The above formulation is applicable to any high-field system and geometry, dc or rf. It provides a general and rigorous formalism to connect the macroscopic experimental findings of breakdown rate dependence to the microscopic arc ignition simulation models presented in sections VI and VII. Furthermore, Eq. (12) is a generalized version of the external coupling circuit model already implemented in ArcPIC and FEMOCS. In fact, the RC circuit implementation of [176, 85] is a special case of Eq. (12).
A systematic simulation of the full coupling of this model with PIC is an ongoing work. However, simple analytic approximations for the impedance functions can be considered for various systems, providing valuable insights and an intuitive understanding of the processes that impede plasma development and result in the empirically observed trends described above. Furthermore, by considering such a simplified approach, as proposed by [145], a direct comparison with empirical data is possible, yielding a very good prediction of the empirically observed limits.
@@
A second simplification that can be done for rf accelerating structures, which gives quite a good agreement with empirical data, is to look at the value of at the single operating frequency of the accelerating structure. At this frequency, the system becomes equivalent to a series resistor circuit, yielding a simple picture of the breakdown impedance, graphically depicted in Fig. 69. In this figure, the straight lines represent the linear dependence of on for three cases, with the slope of the line defining the breakdown impedance . The black curve gives the BD emission characteristic. The crossing point corresponds to the operating field during the breakdown, which [145] proposed as a new empirical limit. In other words, locations in a structure with a higher operating field are expected to have a higher breakdown rate as they maintain a higher flow of electromagnetic power into the breakdown.
The value of needs to be calculated numerically, but simple models can yield analytical formulae. This formula has been derived [145]. By modeling an accelerating structure as a series of coupled LCR circuits, each representing a cell of the rf structure. This model yielded
which was also confirmed by numerical simulations. The dependence on the group velocity, as well as the ratio of surface to accelerating fields, appears consistent with the early empirical predictors of breakdown performance, such as the group velocity or , as the field ratio scales with the circumference.
Beyond this simplification, the value of has been calculated numerically for multiple systems, yielding very promising results that agree with experimental observations. These results can be found in ref. [145]. Here, we present the results for two specific systems, namely the dc large electrode system and the crab cavity.
Figure 70 demonstrates the success of this approach for the crab cavity, in contrast with the picture of Fig. 68. The distribution of the loaded field follows the same trend as the unloaded electric field , thus restoring the prediction that the high-field performance is determined by the power flow rather than the surface electric field value.
Another system where previous high field limits (, etc.) failed to predict the breakdown distribution is the pulsed dc system, as group velocity and are not definable in dc in general. In addition the dc systems show a gap dependence as discussed in Sec. III.5. However, the above approach gave excellent predictions. One of the approximations that is necessary to abandon when studying the dc system is the single frequency. In the dc, there is no reason to choose a certain operating frequency. For that reason, instead of , the indicating limit is given by the RMS value of over a wide range of frequencies. Using this RMS average as a metric, both the positional dependence of breakdown density inside the surface of pulsed dc systems and the gap scaling have been successfully predicted [145].
@@ X.1 Current understanding
In this review, we have summarized a body of experimental, theoretical, and simulation studies that address the fundamental processes of vacuum breakdown. Several notable results significantly advanced the overall understanding of vacuum breakdown. Enabling this advance is a body of consistent and comparable experimental work that allows for a quantitative comparison and combination of data. The experimental work covers high-field and breakdown phenomena over a wider range of conditions and system types than have been before. The work contributes both to fundamental understanding and practical considerations for real-world devices.
@@
One of the most striking theoretical insights is the identification of the central role in the breakdown initiation of the plastic response, i.e., the dislocation dynamics of the electrode material to the stresses caused by surface electromagnetic fields. This insight explains the ordering of field-holding strength of different materials through the different material crystal structures and corresponding energies required for dislocation motion. A formalism based on dislocation dynamics provides a predictive model for the observed dependence of breakdown rate on the applied electric field. This strong dependence has been observed and measured in radio frequency and pulsed dc electrodes. The same formalism also accurately predicts the observed dependence of field-holding on temperature, including the striking increase in field-holding as devices are cooled to cryogenic temperatures. In addition, the dislocation model provides explanations for details of experimentally observed breakdown behavior, for example, the pulse-to-pulse memory effect, which is deduced from the uniform distribution of breakdown timings within pulses, despite a strong dependence of breakdown rate on pulse length.
@@
Analysis of systematic and detailed data from the conditioning and long-term operation of many tens of radio frequency structures and dc electrodes reveals that the underlying conditioning process correlates with the number of pulses, rather than the number of breakdowns. This observation indicates the existence of intrinsic features that are conditioned under applied-field induced cyclic tensile stress. The conditioning of intrinsic features occurs at the highest fields while conditioning of extrinsic features dominate the conditioning of lower electric field level systems, the main domain of previous studies. The way in which the field holding capability of a device evolves in response to the application of fields has practical operational importance but also gives strong insight into the breakdown initiation process. The strong correlation to the number of pulses gives strong evidence that the evolution of the dislocation structure is responsible for the long-term conditioning of high-field devices. This insight into the pulse number dependence has allowed the optimization of conditioning strategies, aided by a Monte Carlo simulation tool that models the conditioning and breakdown process.
@@
Data from testing of multiple device types show a remarkable variation of achievable peak surface field, which turns out to be dependent on device geometry. This means that the macroscopic geometry influences the microscopic breakdown process. The key element in explaining this macroscopic-to-microscopic dependence is the power flow. Simply stated, a certain level of power inflow is necessary to support the breakdown evolution process, in particular while the plasma is forming and ever-increasing currents are emitted from the breakdown site. A series of increasingly comprehensive theoretical models based on the coupling of power flow to breakdowns have been developed. The earliest models considered global power flow, a later one local complex power flow, and finally a full dynamic beam-loading. These successive generations of models quantitatively predict an ever wider range of experiments.
In addition to the dislocation dynamics under the metal surface, the dynamics of surface atoms play an important role in the initiation of the vacuum breakdown. It may lead to sharpening and further growth of surface irregularities whose appearance was triggered by an undersurface dislocation reaction. Surface dynamics is driven by biased diffusion of surface atoms that become partially charged and respond to the presence of a surface gradient. This process is also thermally activated and, hence, slows down at cryogenic temperatures. The process is self-reinforcing; As features grow via biased diffusion, the field gradient increases, leading to stronger diffusion and growth, consistent with the observed critical BD process.
@@
A long-standing mystery of high-field systems is the nature of the features that cause emission enhancement in macroscopic systems. A large correction factor is needed to reconcile the current versus applied field measurements with electron field emission theory, as expressed by the Fowler-Nordheim equation. A widespread assumption is that there are microscopic surface features, commonly called tips, that cause a geometrical field enhancement, which reconciles experiment with theory. Field enhancement factors of 30 are typically needed for the fully conditioned structures described in this report, and factors of 100 to 200 are needed for the dc electrodes. Advanced modeling capabilities have been developed and applied to tips that would give such field enhancements. The models include the response of the tips to the applied field, including heating and cooling processes due to current flow and electron emission, mechanical stresses, and emission of copper atoms, and also include plasma formation and evolution. This modeled the evolution process, including all important physical processes of the breakdown from an initial condition through the observed transition of the vacuum to a conductive medium and collapse of applied fields. The tip simulations provide a simulation basis for the time-evolving impedance used in the power coupling modeling that gives geometrical dependence field holding. Simulations provide insight into different aspects of breakdown sites, including, for example, constraints on the nature of the tips and thermal coupling to the base.
Reduction of work function due to the application of an electric field was also considered as a mechanism to enhance field emission. It was found to be unlikely to expect a significant effect as a result of electron density redistribution around the surface atoms exposed to the external electric field; however, some experimental work reports an observed reduction in work function. A variation of the work function value is observed both theoretically and experimentally for different surface facets; however, further theoretical studies are needed to investigate whether the bulk defects developed under the applied pulsed electric field produce a sufficient variation in the work function.
@@ X.2 Prospects and open questions
This report describes a coherent theoretical foundation for a large body of mature experimental work. There are, inevitably, several open questions some of which are the subject of ongoing efforts, in particular in connecting remaining important theoretical constructs with direct experimental observations. In addition there are efforts underway to apply the insights and methodologies described here to a wider range devices, with a correspondingly wider range of operating conditions and parameters. This expansion will hopefully help these applications, but will also provide new data acquired in a way that can compared qualitatively with those presented here. This will provide further testing of ideas and models and help improve our understanding of breakdown.
@@
One of these ongoing research focus areas is to identify exactly what is changing as a device conditions, through direct observations of how a conditioned surface and subsurface differs from an unconditioned one. As we have seen, the dislocation model described in Sec. IV predicts the field and temperature dependence of breakdown rate. An extension of the role of dislocations is that their structure or population changes as conditioning proceeds, in analogy to work hardening of metals. One of the primary experimental lines of this effort is to perform microscopic imaging of devices at various conditioning stages and, in complement, as a function of position in spatially varying surface fields. The primary experimental tool is electron microscopy, for example with FIB sample preparation followed by TEM analysis.
@@
Another major open question is the nature of the enhanced emission sites, and how they populate high-field surfaces. Data from field emission measurements from all macroscopic devices require significant emission enhancement factors as we have seen in Secs. II and VI. But at the time of this writing, no sufficiently sharp tip-like feature has ever been found in many hundreds of SEM images taken in the course of the experiments described in this report nor is anyone able to predict the emission enhancement factor independently of a field emission measurement. Imaging of of the spatial distribution of field emission is being carried out by upgrading the breakdown positioning cameras of pulsed dc systems described in II to higher sensitivity. The upgraded system will allow the question of whether breakdowns actually preferentially occur where stable field emitters are observed, or not, to be answered.
Examples of devices where the experimental methodology described here is now being applied include radio frequency quadrupoles, electrostatic separators and muon capture cavities, all devices used in accelerators but much larger than those described here and operating at lower peak fields. Additional effects reduce field holding; H- irradiation in radio frequency cavities [163], irradiation in electrostatic separators and immersion in multi-tesla external magnetic fields for muon capture cavities. The expanded parameter space and the role of additional breakdown-inducing effects will help us refine our understanding.
Finally, there is a clear role for an experimental set-up that combines a high-field device with in-vacuum diagnostic systems. This could be, for example, a pulsed dc system combined with in-vacuum manipulators to transfer electrodes to a scanning electron microscope, ideally combined with a focused ion beam system or a scanning tunneling microscope. Such a setup could help identify the origin of field emission sites or identify evidence of the link between plastic activity and field emission leading to BD processes. Direct observation of these effects is made complicated when samples are exposed to air between exposure to high-fields and analysis. Initial attempts at in-vacuum experiments have been made, but technical challenges have not allowed robust and conclusive results [129]. We are hopeful that similar work, if pursued in the future, will provide a direct observational link.
Appendix A Glossary
| Term | Definition | Abbr. |
| Vacuum arc | A plasma appearing near the metal surface under vacuum conditions after high electric field exposure | — |
| Vacuum Breakdown | Transition from an insulating to a conducting state via a localized plasma above the surface | VBD |
| Breakdown Rate | number of pulses with breakdown divided by the total number of pulses in a measurement window | BDR |
| Conditioning | controlled process of gradually increasing field and/or pulse length to improve field-holding capability | — |
| Field-holding capability | ability to sustain a given surface electric field without breakdown | — |
| Field (electron) emission | electron emission by quantum tunneling from a metal surface under strong electric fields | FE |
| Dark current | Current caused by field emitted electrons | — |
| Maxwell stress | Tensile stress acting on a conductor’s surface due to the applied electric field | — |
| Biased diffusion | Diffusion of surface atoms with uneven jump rate didstribution, resulting in atomic migration toward regions of higher electric field | — |
| Vacuum bursts | sudden pressure spikes in a vacuum system | — |
| Large Electrode System | A pulsed dc test system with two large, flat electrodes separated by a precision spacer | LES |
| Anode tip system | pulsed dc setup with a shaped anode tip facing a flat cathode at micrometer-scale gaps | — |
| Faraday cup | current detector that collects charged particles along a beam axis | — |
| Modified Poyinting vector | local power-flow metric characterizing power available to initiate an arc | |
| Group velocity () | speed at which rf power propagates through a traveling-wave structure | |
| Accelerating gradient | on-axis average field experienced by the beam in an accelerating structure | — |
| Pulse length | duration of the applied rf (or dc) high-voltage pulse | |
| Thermal runaway | positive-feedback heating of an emitting nano-tip leading to melting, evaporation, and enhanced emission | TR |
| Plasma sheath | near-surface region where potential drops by 20 V (“burning voltage” for Cu arcs), accelerating ions into the cathode | — |
| Breakdown impedance | effective local impedance coupling the microscopic breakdown current to the macroscopic electromagnetic environment | Z |